OpenStudy (anonymous):

Find the first 5 terms of the sequence: a1 = 500, an = (an-1)/5.

OpenStudy (amistre64):

\(\color{red}{\small\text{since they give you the formula, just use it :)}}\)

OpenStudy (amistre64):

a{2} = (500-2)/5

OpenStudy (amistre64):

a{3} = a{2}-1/5 got a spurious typo in the first try

OpenStudy (anonymous):

500, 100, 20, 5, 1?

OpenStudy (amistre64):

500 500/5 = 100 100/5 = 20 ; yep

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

\(\color{green}{\small\text{good job :)}}\)

OpenStudy (anonymous):

thx :)